# The Learning Dynamics of Subword Segmentation for Morphologically Diverse Languages

### v1

@@ 1 Introduction
Subword tokenisers like BPE (sennrich-etal-2016-neural) and ULM (kudo-2018-subword) are applied during preprocessing, after which they remain unchanged during language model (LM) pretraining and finetuning. If tokenisers produce subwords that are ineffective units for language modelling, this cannot be fixed during training. This is especially problematic for settings in which models are sensitive to tokenisation decisions, such as low-resource tasks (zhu-etal-2019-importance; wang-etal-2021-multi-view) and morphologically rich languages (zhu-etal-2019-systematic; klein-tsarfaty-2020-getting).
@@
Ideally, subword segmentation should be learned during end-to-end training. Previous works achieve this by jointly optimising segmentation and model parameters (kreutzer-sokolov-2018-learning; he-etal-2020-dynamic; meyer-buys-2022-subword). In this paper, we use this paradigm to study the learning dynamics of subword segmentation. When an LM is able to learn subword segmentation, how do its subword boundaries evolve over the course of pretraining and finetuning? The question has practical implications, as it reveals how tokenisation requirements vary across different stages of training. There is no guarantee that subword units that are optimal for the early epochs of pretraining would also be suitable for later stages of pretraining, or meet the demands of subsequent task-specific finetuning.
Subword learning dynamics may also depend on linguistic properties, particularly morphology. We conduct experiments on three languages spanning a range of morphological complexity: Setswana, English, and isiXhosa. Setswana and isiXhosa are South African languages that represent opposite ends of the morpho-orthographic spectrum. Both are classified as agglutinative, but they differ in how orthographic (written) words are constructed (taljard-bosch-2006-comparison). IsiXhosa is conjunctive: morphemes are concatenated into long word forms, whose meaning depends on morphological decomposition. Setswana is disjunctive: morphemes are written as space-separated words and many words consist of single morphemes. English, with its limited inflectional morphology, serves as a typological middle ground. The examples in Table 1 illustrate the contrasting subword structures.
Our study requires a model capable of dynamically learning subword segmentation during pretraining and finetuning. For this purpose we extend the subword segmental language model (SSLM) (meyer-buys-2022-subword), which jointly optimises subword segmentation and LM training. We develop a Transformer-based variant, T-SSLM, incorporating learnable subword segmentation into modern LM practices like pretraining, task-specific finetuning, and text generation.
@@
We pretrain T-SSLMs for Setswana, English, and isiXhosa, tracking learned subwords across checkpoints from three linguistic perspectives. Morphological alignment measures how closely subwords align with morphological boundaries. Productivity (gutierrez-vasques-etal-2023-languages) measures the generative capacity of subwords, while Fertility (acs-2021-exploring) estimates lexical coverage. Across all three criteria, isiXhosa exhibits the greatest instability, while Setswana converges earliest. We identify distinctive phases of subword learning, characterised by clear changes and trends in our linguistic metrics. Our results show that different stages of training and different languages necessitate characteristically different subword units.
@@ 2.1 Learning dynamics of language models
Previous research tracked the evolution of syntactic and semantic knowledge during training. Some studies track internal representation change (saphra-lopez-2019-understanding; chiang-etal-2020-pretrained; liu-etal-2021-probing-across; muller-eberstein-etal-2023-subspace), while others track performance on grammatical tasks (choshen-etal-2022-grammar; evanson-etal-2023-language). In all these studies (as in most LM research) tokenisation is fixed, so there is no subword learning trajectory to study. gutierrez-vasques-etal-2023-languages analyse how BPE token properties change over successive merges; however, BPE is still fixed after preprocessing, so the emergence of subwords in LM training is not studied. That requires considering approaches that incorporate subword learning into training.
@@ 2.2 Unifying training and tokenisation
Previous works unify training and subword learning by marginalising over different candidate tokenisations during training. This has been used to learn subwords for translation (kreutzer-sokolov-2018-learning; he-etal-2020-dynamic; meyer-buys-2023-subword) and unsupervised Chinese word segmentation (sun-deng-2018-unsupervised; kawakami-etal-2019-learning; downey-etal-2022-masked). meyer-buys-2022-subword adapt this technique to propose subword segmental language modelling (SSLM), an LSTM-based LM architecture that dynamically optimises subword segmentation. They train small-scale SSLMs for low-resource, agglutinative languages. They do not conduct finetuning or downstream evaluation, but perplexity-based evaluation demonstrates the potential of SSLM over tokenisation-based LMs.
@@ 3 Transformer SSLM
We leverage the technique underlying the SSLM of meyer-buys-2022-subword – learning subword units during LM training by marginalising over potential tokenisations – as an opportunity to study subword learning dynamics. The original SSLM architecture is based on a shallow LSTM. To align SSLM with current best practices for pretraining, task-finetuning, and open-ended text generation, we develop a Transformer version of the model. We propose T-SSLM, a Transformer-based adaptation of SSLM that brings the model up to date with modern LM architectures.
From a modelling perspective, our T-SSLM is a conceptually straightforward but technically non-trivial extension of the original SSLM. We do not position our architectural changes as the main contribution of this paper. Our primary research goal is empirical analysis: to study subword learning dynamics across different stages of training. We develop T-SSLM as a new tool for studying subword segmentation as a learnable component of language modelling. We will release our implementation of T-SSLM upon publication. [^1] https://anonymous.4open.science/r/transformer-sslm-E154
In this section, we present a technical overview of how T-SSLM dynamically learns subword segmentation during pretraining and finetuning. For a comprehensive introduction to the underlying SSLM framework, we refer readers to meyer-buys-2022-subword. Here we summarise the method behind the original SSLM, outline how we adapt it to Transformer-based modelling, and explain how we use T-SSLM to track subword learning.
@@ 3.1 SSLM background
The number of possible subword segmentations of a document $D$ grows exponentially with its length, so the marginalisation in Eq. 2 becomes intractable. meyer-buys-2022-subword propose a dynamic programming algorithm to compute Eq. 2 efficiently. Suppose document $D$ consists of characters $\mathbf{c}=c_{1},c_{2},\ldots,c_{|D|}$. The algorithm computes the marginalised probability up to the $k^{\mathrm{th}}$ character as $p(D_{1:k})=\alpha_{k}$, where $\alpha_{0}=1$ and $\alpha_{k}$ are forward scores computed as
@@ 3.2 T-SSLM: adapting SSLM for pretraining and finetuning
meyer-buys-2022-subword parameterise the SSLM described above with an LSTM (hochreiter-schmidhuber-1997-long), so the subword probabilities in Eq. 3 are conditioned on LSTM encodings. We propose T-SSLM, a new variant of SSLM parameterised by a Transformer (vaswani-etal-2017-attention). This enables larger-scale pretraining and aligns SSLM with modern LM architectures.
To construct a neural LM that can compute $p_{\theta}(D,T)$ for any candidate segmentation $T$, meyer-buys-2022-subword introduce several architectural innovations, which we replicate in T-SSLM. One of these innovations is to encode the preceding text $t_{<i}$ as an untokenised character sequence $c_{<j}$ to achieve computationally feasible conditioning on the tokenisation history. In T-SSLM, we use a character-level Transformer for this encoding instead of an LSTM. This Transformer forms the backbone of our architecture: the subword probabilities computed in Eq. 3 are conditioned on its learned output representations.
@@ 4.1 Pretraining
T-SSLM’s time complexity (detailed in §8.1) makes it computationally challenging to scale the pretraining data size to match state-of-the-art LMs. However, we view the technique as an opportunity for analysis: a unique lens through which to analyse the role of subword segmentation in LM training. For pretraining, we prioritise high-quality corpora over scale to ensure that models are trained on clean, representative language input. Our corpus sizes are listed in Table 2. For isiXhosa we use the 13m-word WURA oladipo-etal-2023-better corpus, created by filtering mC4 xue-etal-2021-mt5 to discard low-quality data. For the lower-resourced Setswana we use the 4m-word PuoData corpus (marivate2023puoberta). For English we use the 10m-word BabyLM corpus (warstadt-etal-2023-findings). To track learning dynamics, we store model checkpoints at regular pretraining intervals. After pretraining, we analyse the subwords of nine checkpoints spread evenly across pretraining: one checkpoint after the first epoch, one after the final epoch, and seven additional checkpoints spaced at regular intervals (1/8ths of total pretraining).
@@ 4.2 Finetuning
For data-to-text we use the T2X dataset (meyer-buys-2024-triples). It contains triples (subject, relation, object) paired with descriptive isiXhosa sentences e.g. (South Africa, leader, Cyril Ramaphosa) $\rightarrow$ “uCyril Ramaphosa yinkokheli yoMzantsi Afrika” (“Cyril Ramaphosa is the leader of South Africa”). For headline generation from the body of a news article we use MasakhaNEWS (adelani-etal-2023-masakhanews), which contains news articles paired with headlines. This is a much harder task than data-to-text, especially considering the smaller finetuning dataset (see Table 2).
@@ 4.3 Baselines
For each language, we pretrain and finetune five LMs with fixed subword tokenisation (hyperparameters are reported in Appendix C). We use the same datasets as our T-SSLMs to ensure a fair comparison. Three baselines use well-established tokenisers: (1) BPE (sennrich-etal-2016-neural), (2) ULM (kudo-2018-subword), and (3) BPE-dropout (provilkov-etal-2020-bpe). Additionally, we use (4) character-level segmentation and (5) byte-based modelling, both of which have been shown to outperform subword-based LMs on low-resource languages (edman-etal-2022-subword; edman-etal-2024-character; adelani-etal-2022-thousand).
@@ 4.4.1 Morphological alignment
Our models do not have access to morphological annotations, but we investigate to what extent they “discover” morphemes as subword units by tracking the amount of overlap between learned subwords and morphemes. For Setswana and isiXhosa we use the NCHLT test corpus (eiselen-puttkammer-2014-developing), which contains 5,000 morphologically annotated words for each language. For English we use the 2022 SIGMORPHON dataset (batsuren-etal-2022-sigmorphon), which contains 18,000 morphologically segmented words. We compare the subwords learned by our models to the gold-standard morphological segmentations. To measure morphological alignment, we compute precision, recall, and F1 for morphological boundary identification.
@@ 4.4.2 Productivity and Idiosyncrasy
Morphological productivity is the capacity of a morpheme to combine with other morphemes to form new words (Bybee_2001). Productive morphemes occur in many word types, often forming regular morphological patterns. For example, in isiXhosa prefixes can be attached to stems to indicate plurality e.g. “abahlobo” (friends) and “abazali” (parents). In contrast, idiosyncratic morphemes are contained in few word types, but these types occur frequently. For example, the isiXhosa stem “–ntu” (human) will not occur in as many unique words as prefixal morphemes, but the words in which “–ntu” occurs are commonly used e.g. “umntu” (person) and “abantu” (people).
We quantify productivity and idiosyncrasy as proposed by gutierrez-vasques-etal-2023-languages,
@@ 4.4.3 Fertility
The fertility of a tokeniser is the average number of subwords per segmented word (acs-2021-exploring). It is a measure of lexical coverage – the extent to which the words of a language are included in the tokeniser vocabulary. A fertility of 1 indicates that every word in a corpus is included in the vocabulary (no segmented words), while a higher fertility corresponds to finer-grained segmentation and lower lexical coverage. We also compute fertility based on the segmentation of our LM validation corpora.
@@ 5 Subword Learning Dynamics
Figure 3 shows similarly rapid changes in subword productivity. For English and isiXhosa, productivity and idiosyncrasy decrease at a constant rate. gutierrez-vasques-etal-2023-languages found a similar pattern in early BPE merges. Highly productive, high-frequency subwords are identified at the outset, after which the model incorporates progressively less common subwords. This leads to decreased productivity and idiosyncrasy. For Setswana, the learning trajectory is different. Productivity also decreases, but idiosyncrasy remains stable. This also agrees with gutierrez-vasques-etal-2023-languages, who showed that BPE tokens for morphological simple languages have low productivity but high idiosyncrasy. Many Setswana subwords are near-whole words, so they occur in few types, but the types are not rare (i.e. high idiosyncrasy).
@@
During finetuning, fertility increases for all three languages (Figure 1). Previous work showed finer-grained units, such as characters or bytes, are beneficial for low-resource languages with small finetuning datasets (edman-etal-2022-subword; edman-etal-2024-character; adelani-etal-2022-thousand). More aggressive segmentation may also suit the data-to-text dataset, since it contains many named entities (e.g. “Akeen Dent” in Table 4). Names cannot be segmented into semantically coherent subwords, since they do not consist of morphemes. It is reasonable to model names as character sequences, which is what we observe after finetuning. We present examples of this change from pretraining to finetuning in Table 4.
@@ 6 Downstream Text Generation
We finetune T-SSLM and five tokenisation-based baseline LMs (see §4.3) on two isiXhosa text generation tasks (data-to-text and headline generation, see §4.2). We evaluate performance with BLEU (papineni-etal-2002-bleu) and chrF (popovic-2015-chrf), which quantifies subword-level overlap. Additionally we measure the proportion of test examples that produce degenerative text (incoherent or repetitive language). A generation is considered degenerative if a word is repeated at least three times (word-level repetition) or if a word is longer than 30 characters (subword-level repetition).
The results in Table 5 shows that T-SSLM outperforms tokenisation-based models across reference-based metrics, with substantial BLEU gains. It overcomes text degeneration to a considerable degree for in-language (isiXhosa) pretraining and finetuning, producing much less repetition (Table 7 in the appendix shows examples of generated text). Text degeneration is a challenging problem in low-resource tasks, as evidenced by the high proportions of degenerative examples produced by baselines. T-SSLM does not match the performance of massively multilingual LMs (meyer-etal-2024-nglueni), but our aim is to isolate the effect of task-adaptable subwords over fixed tokenisation. We therefore focus on comparing SSLM to tokenisation-based LMs of the same size trained on the same datasets. In this comparison our results demonstrate the benefit of learnable subwords in data-scarce settings.

### v2

@@ 1 Introduction
Subword tokenisers like BPE (Sennrich et al., 2016) and ULM (Kudo, 2018) are applied during preprocessing, after which they remain unchanged during language model (LM) pretraining and finetuning. If tokenisers produce subwords that are ineffective units for language modelling, this cannot be fixed during training. This is especially problematic for settings in which models are sensitive to tokenisation decisions, such as low-resource tasks (Zhu et al., 2019a; Wang et al., 2021) and morphologically rich languages (Zhu et al., 2019b; Klein and Tsarfaty, 2020).
@@
Ideally, subword segmentation should be learned during end-to-end training. Previous works achieve this by jointly optimising segmentation and model parameters (Kreutzer and Sokolov, 2018; He et al., 2020; Meyer and Buys, 2022). In this paper, we use this paradigm to study the learning dynamics of subword segmentation. When an LM is able to learn subword segmentation, how do its subword boundaries evolve over the course of pretraining and finetuning? The question has practical implications, as it reveals how tokenisation requirements vary across different stages of training. There is no guarantee that subword units that are optimal for the early epochs of pretraining would also be suitable for later stages of pretraining, or meet the demands of subsequent task-specific finetuning.
Subword learning dynamics may also depend on linguistic properties, particularly morphology. We conduct experiments on three languages spanning a range of morphological complexity: Setswana, English, and isiXhosa. Setswana and isiXhosa are South African languages that represent opposite ends of the morpho-orthographic spectrum. Both are classified as agglutinative, but they differ in how orthographic (written) words are constructed (Taljard and Bosch, 2006). IsiXhosa is conjunctive: morphemes are concatenated into long word forms, whose meaning depends on morphological decomposition. Setswana is disjunctive: morphemes are written as space-separated words and many words consist of single morphemes. English, with its limited inflectional morphology, serves as a typological middle ground. The examples in Table 1 illustrate the contrasting subword structures.
Our study requires a model capable of dynamically learning subword segmentation during pretraining and finetuning. For this purpose we extend the subword segmental language model (SSLM) (Meyer and Buys, 2022), which jointly optimises subword segmentation and LM training. We develop a Transformer-based variant, T-SSLM, incorporating learnable subword segmentation into modern LM practices like pretraining, task-specific finetuning, and text generation.
@@
We pretrain T-SSLMs for Setswana, English, and isiXhosa, tracking learned subwords across checkpoints from three linguistic perspectives. Morphological alignment measures how closely subwords align with morphological boundaries. Productivity (Gutierrez-Vasques et al., 2023) measures the generative capacity of subwords, while Fertility (Ács, 2019) estimates lexical coverage. Across all three criteria, isiXhosa exhibits the greatest instability, while Setswana converges earliest. We identify distinctive phases of subword learning, characterised by clear changes and trends in our linguistic metrics. Our results show that different stages of training and different languages necessitate characteristically different subword units.
@@ 2.1 Learning dynamics of language models
Previous research tracked the evolution of syntactic and semantic knowledge during training. Some studies track internal representation change (Saphra and Lopez, 2019; Chiang et al., 2020; Liu et al., 2021; Müller-Eberstein et al., 2023), while others track performance on grammatical tasks (Choshen et al., 2022; Evanson et al., 2023). In all these studies (as in most LM research) tokenisation is fixed, so there is no subword learning trajectory to study. Gutierrez-Vasques et al. (2023) analyse how BPE token properties change over successive merges; however, BPE is still fixed after preprocessing, so the emergence of subwords in LM training is not studied. That requires considering approaches that incorporate subword learning into training.
@@ 2.2 Unifying training and tokenisation
Previous works unify training and subword learning by marginalising over different candidate tokenisations during training. This has been used to learn subwords for translation (Kreutzer and Sokolov, 2018; He et al., 2020; Meyer and Buys, 2023) and unsupervised Chinese word segmentation (Sun and Deng, 2018; Kawakami et al., 2019; Downey et al., 2022). Meyer and Buys (2022) adapt this technique to propose subword segmental language modelling (SSLM), an LSTM-based LM architecture that dynamically optimises subword segmentation. They train small-scale SSLMs for low-resource, agglutinative languages. They do not conduct finetuning or downstream evaluation, but perplexity-based evaluation demonstrates the potential of SSLM over tokenisation-based LMs.
@@ 3 Transformer SSLM
We leverage the technique underlying the SSLM of Meyer and Buys (2022) – learning subword units during LM training by marginalising over potential tokenisations – as an opportunity to study subword learning dynamics. The original SSLM architecture is based on a shallow LSTM. To align SSLM with current best practices for pretraining, task-finetuning, and open-ended text generation, we develop a Transformer version of the model. We propose T-SSLM, a Transformer-based adaptation of SSLM that brings the model up to date with modern LM architectures.
From a modelling perspective, our T-SSLM is a conceptually straightforward but technically non-trivial extension of the original SSLM. We do not position our architectural changes as the main contribution of this paper. Our primary research goal is empirical analysis: to study subword learning dynamics across different stages of training. We develop T-SSLM as a new tool for studying subword segmentation as a learnable component of language modelling. Our implementation of T-SSLM is publicly available. [^1] https://github.com/francois-meyer/transformer-sslm
In this section, we present a technical overview of how T-SSLM dynamically learns subword segmentation during pretraining and finetuning. For a comprehensive introduction to the underlying SSLM framework, we refer readers to Meyer and Buys (2022). Here we summarise the method behind the original SSLM, outline how we adapt it to Transformer-based modelling, and explain how we use T-SSLM to track subword learning.
@@ 3.1 SSLM background
The number of possible subword segmentations of a document $D$ grows exponentially with its length, so the marginalisation in Eq. 2 becomes intractable. Meyer and Buys (2022) propose a dynamic programming algorithm to compute Eq. 2 efficiently. Suppose document $D$ consists of characters $\mathbf{c}=c_{1},c_{2},\ldots,c_{|D|}$. The algorithm computes the marginalised probability up to the $k^{\mathrm{th}}$ character as $p(D_{1:k})=\alpha_{k}$, where $\alpha_{0}=1$ and $\alpha_{k}$ are forward scores computed as
@@ 3.2 T-SSLM: adapting SSLM for pretraining and finetuning
Meyer and Buys (2022) parameterise the SSLM described above with an LSTM (Hochreiter and Schmidhuber, 1997), so the subword probabilities in Eq. 3 are conditioned on LSTM encodings. We propose T-SSLM, a new variant of SSLM parameterised by a Transformer (Vaswani et al., 2017). This enables larger-scale pretraining and aligns SSLM with modern LM architectures.
To construct a neural LM that can compute $p_{\theta}(D,T)$ for any candidate segmentation $T$, Meyer and Buys (2022) introduce several architectural innovations, which we replicate in T-SSLM. One of these innovations is to encode the preceding text $t_{<i}$ as an untokenised character sequence $c_{<j}$ to achieve computationally feasible conditioning on the tokenisation history. In T-SSLM, we use a character-level Transformer for this encoding instead of an LSTM. This Transformer forms the backbone of our architecture: the subword probabilities computed in Eq. 3 are conditioned on its learned output representations.
@@ 4.1 Pretraining
T-SSLM’s time complexity (detailed in §8.1) makes it computationally challenging to scale the pretraining data size to match state-of-the-art LMs. However, we view the technique as an opportunity for analysis: a unique lens through which to analyse the role of subword segmentation in LM training. For pretraining, we prioritise high-quality corpora over scale to ensure that models are trained on clean, representative language input. Our corpus sizes are listed in Table 2. For isiXhosa we use the 13m-word WURA Oladipo et al. (2023) corpus, created by filtering mC4 Xue et al. (2021) to discard low-quality data. For the lower-resourced Setswana we use the 4m-word PuoData corpus (Marivate et al., 2023). For English we use the 10m-word BabyLM corpus (Warstadt et al., 2023). To track learning dynamics, we store model checkpoints at regular pretraining intervals. After pretraining, we analyse the subwords of nine checkpoints spread evenly across pretraining: one checkpoint after the first epoch, one after the final epoch, and seven additional checkpoints spaced at regular intervals (1/8ths of total pretraining).
@@ 4.2 Finetuning
For data-to-text we use the T2X dataset (Meyer and Buys, 2024). It contains triples (subject, relation, object) paired with descriptive isiXhosa sentences e.g. (South Africa, leader, Cyril Ramaphosa) $\rightarrow$ “uCyril Ramaphosa yinkokheli yoMzantsi Afrika” (“Cyril Ramaphosa is the leader of South Africa”). For headline generation from the body of a news article we use MasakhaNEWS (Adelani et al., 2023), which contains news articles paired with headlines. This is a much harder task than data-to-text, especially considering the smaller finetuning dataset (see Table 2).
@@ 4.3 Baselines
For each language, we pretrain and finetune five LMs with fixed subword tokenisation (hyperparameters are reported in Appendix C). We use the same datasets as our T-SSLMs to ensure a fair comparison. Three baselines use well-established tokenisers: (1) BPE (Sennrich et al., 2016), (2) ULM (Kudo, 2018), and (3) BPE-dropout (Provilkov et al., 2020). Additionally, we use (4) character-level segmentation and (5) byte-based modelling, both of which have been shown to outperform subword-based LMs on low-resource languages (Edman et al., 2022, 2024; Adelani et al., 2022).
@@ 4.4.1 Morphological alignment
Our models do not have access to morphological annotations, but we investigate to what extent they “discover” morphemes as subword units by tracking the amount of overlap between learned subwords and morphemes. For Setswana and isiXhosa we use the NCHLT test corpus (Eiselen and Puttkammer, 2014), which contains 5,000 morphologically annotated words for each language. For English we use the 2022 SIGMORPHON dataset (Batsuren et al., 2022), which contains 18,000 morphologically segmented words. We compare the subwords learned by our models to the gold-standard morphological segmentations. To measure morphological alignment, we compute precision, recall, and F1 for morphological boundary identification.
@@ 4.4.2 Productivity and Idiosyncrasy
Morphological productivity is the capacity of a morpheme to combine with other morphemes to form new words (Bybee, 2001). Productive morphemes occur in many word types, often forming regular morphological patterns. For example, in isiXhosa prefixes can be attached to stems to indicate plurality e.g. “abahlobo” (friends) and “abazali” (parents). In contrast, idiosyncratic morphemes are contained in few word types, but these types occur frequently. For example, the isiXhosa stem “–ntu” (human) will not occur in as many unique words as prefixal morphemes, but the words in which “–ntu” occurs are commonly used e.g. “umntu” (person) and “abantu” (people).
We quantify productivity and idiosyncrasy as proposed by Gutierrez-Vasques et al. (2023),
@@ 4.4.3 Fertility
The fertility of a tokeniser is the average number of subwords per segmented word (Ács, 2019). It is a measure of lexical coverage – the extent to which the words of a language are included in the tokeniser vocabulary. A fertility of 1 indicates that every word in a corpus is included in the vocabulary (no segmented words), while a higher fertility corresponds to finer-grained segmentation and lower lexical coverage. We also compute fertility based on the segmentation of our LM validation corpora.
@@ 5 Subword Learning Dynamics
Figure 3 shows similarly rapid changes in subword productivity. For English and isiXhosa, productivity and idiosyncrasy decrease at a constant rate. Gutierrez-Vasques et al. (2023) found a similar pattern in early BPE merges. Highly productive, high-frequency subwords are identified at the outset, after which the model incorporates progressively less common subwords. This leads to decreased productivity and idiosyncrasy. For Setswana, the learning trajectory is different. Productivity also decreases, but idiosyncrasy remains stable. This also agrees with Gutierrez-Vasques et al. (2023), who showed that BPE tokens for morphological simple languages have low productivity but high idiosyncrasy. Many Setswana subwords are near-whole words, so they occur in few types, but the types are not rare (i.e. high idiosyncrasy).
@@
During finetuning, fertility increases for all three languages (Figure 1). Previous work showed finer-grained units, such as characters or bytes, are beneficial for low-resource languages with small finetuning datasets (Edman et al., 2022, 2024; Adelani et al., 2022). More aggressive segmentation may also suit the data-to-text dataset, since it contains many named entities (e.g. “Akeen Dent” in Table 4). Names cannot be segmented into semantically coherent subwords, since they do not consist of morphemes. It is reasonable to model names as character sequences, which is what we observe after finetuning. We present examples of this change from pretraining to finetuning in Table 4.
@@ 6 Downstream Text Generation
We finetune T-SSLM and five tokenisation-based baseline LMs (see §4.3) on two isiXhosa text generation tasks (data-to-text and headline generation, see §4.2). We evaluate performance with BLEU (Papineni et al., 2002) and chrF (Popović, 2015), which quantifies subword-level overlap. Additionally we measure the proportion of test examples that produce degenerative text (incoherent or repetitive language). A generation is considered degenerative if a word is repeated at least three times (word-level repetition) or if a word is longer than 30 characters (subword-level repetition).
The results in Table 5 shows that T-SSLM outperforms tokenisation-based models across reference-based metrics, with substantial BLEU gains. It overcomes text degeneration to a considerable degree for in-language (isiXhosa) pretraining and finetuning, producing much less repetition (Table 7 in the appendix shows examples of generated text). Text degeneration is a challenging problem in low-resource tasks, as evidenced by the high proportions of degenerative examples produced by baselines. T-SSLM does not match the performance of massively multilingual LMs (Meyer et al., 2024), but our aim is to isolate the effect of task-adaptable subwords over fixed tokenisation. We therefore focus on comparing SSLM to tokenisation-based LMs of the same size trained on the same datasets. In this comparison our results demonstrate the benefit of learnable subwords in data-scarce settings.